OpenStudy (anonymous):

Solve by the linear combination method (with or without multiplication). 4g + 5h = 4 8g – 10h = 4

OpenStudy (texaschic101):

Is linear combination the same as elimination ??

OpenStudy (anonymous):

I think so but I'm not sure.

OpenStudy (btaylor):

In order to combine, you have to multiply the top equation by 2 so you get 8g + 10h = 8 8g - 10h = 4 --------------- 16g = 12 g = 3/4 --> plug it in to solve for h 8(3/4)-10h = 4 6-10h=4 10h=2 h=1/5 Your result: g=3/4, h=1/5

OpenStudy (texaschic101):

I agree with BTaylor....that is exactly the way I would do it :)